OpenStudy (anonymous):

Jeffrey needs to spend at least 11 more hours studying for his college entrance exam. Tomorrow he has three more hours available than today. The next day he has half the time he does today. How many hours should he study each day?”

OpenStudy (anonymous):

Translate the following situation into an inequality. Do not solve.

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

4 hours the first 5 the second 2 the third

OpenStudy (anonymous):

can you tranlate that into an inequality, please?

OpenStudy (anonymous):

2<X<5 - - those are less than or equal to and greater than or equal to

OpenStudy (anonymous):

thats the final answer

OpenStudy (a_clan):

\[x+(x+3)+(x/2) \ge 11\] x= time available today (x+3) = time available tomorrow (x/2)= time available the next day